### v1

# Hybrid Beamforming with Widely-spaced-array for Multi-user Cross-Near-and-Far-Field Communications

Abstract
With multi-GHz bandwidth, Terahertz (THz) beamforming has drawn increasing attention in the sixth generation (6G) and beyond communications. Existing beamforming designs mainly focus on a compact antenna array where typical communication occurs in the far-field. However, in dense multi-user scenarios, only relying on far-field angle domain fails to distinguish users at similar angles. Therefore, a multi-user widely-spaced array (MU-WSA) is exploited in this paper, which enlarges the near-field region to introduce the additional distance domain, leading to a new paradigm of cross-near-and-far-field (CNFF) communication. Under this paradigm, the CNFF channel model is investigated, based on which the subarray spacing $d_{s}$ and the number of subarrays $K$ in MU-WSA are optimized to maximize the channel capacity. Then, in sub-connected systems, an alternating optimization (AO) beamforming algorithm is proposed to deal with the special block-diagonal format of the analog precoder. For fully-connected systems, a low-complexity steering-vector reconstruction (SVR)-based algorithm is proposed by constructing specialized steering vectors of MU-WSA. Numerical evaluations show that due to distance domain resolutions, the MU-WSA can improve the SE by over $60$at a power of $20$dBm compared to the compact array. Additionally, the proposed AO algorithm in the SC system can achieve over 80% of the sum (SE) of the FC system while reducing the number of phase shifters by $K^{2}$, thereby lowering power consumption. The SVR algorithm in the FC system can achieve over 95% of the upper bound of SE but takes only 10% of the running time of the singular vector decomposition (SVD)-based algorithms.
Index Terms:
Terahertz communications, Multi-user hybrid beamforming, Cross near-and-far-field, Widely-spaced array.
I Introduction
Terahertz (THz) communication has been regarded as an enabling technology for 6G and beyond wireless networks thanks to its ultra-broad continuous bandwidth, ranging from $0.1$ to $10$ THz [1]. However, due to much higher carrier frequency, the severe path loss at the THz band significantly limits the transmission distance and the coverage range [2]. Fortunately, by exploiting the sub-millimeter wavelength, the ultra-massive multiple-input multiple-output (UM-MIMO) antenna array can be utilized as a potential solution to achieve high array gain and compensate for the propagation loss [3]. However, with traditional fully-digital (FD) beamforming architectures where each antenna is equipped with one dedicated RF chain, the huge antenna array brings high hardware complexity and power consumption [4]. To tackle this problem, hybrid beamforming with fewer RF chains is proposed as an appealing technology, which is able to achieve near-optimal performance with acceptable hardware complexity [5].
Most existing hybrid beamforming schemes consider far-field transmission where the electromagnetic (EM) wavefront is approximated as a plane such that the planar wave model (PWM) is utilized to characterize the channel [6]. Under this assumption, the fully-connected (FC) and the sub-connected (SC) architectures are widely-explored. In FC where each RF chain connects to all the antennas, near-optimal spectral efficiency (SE) can be obtained [7, 8, 9, 10]. While in SC where each RF chain only connects to a subset of antennas, the power consumption and the hardware complexity can be reduced at the cost of SE [11, 12, 13].
However, only considering the far-field PWM imposes two limitations for THz multi-user communications. On one hand, the number of transmit data streams for each user is upper-bounded by the number of multipath components, which could be very limited at the THz band, e.g., less than 5 [14, 15, 16]. On the other hand, as far-field beamforming can only provide angle domain resolution, it cannot distinguish users located at the same or similar angles [17, 18]. Therefore, far-field beamforming lacks the ability to support ubiquitous connectivity in the 6G vision, e.g. $10^{7}/\text{km}^{2}$ [19]. To tackle these problems, near-field beamforming has been proposed to provide additional spatial degrees of freedom (SDoF) due to the non-negligible spherical wavefront in near-field [20, 21]. Leveraging this property, the authors in [20] proposed different beam focusing schemes for multi-users in the fully digital, hybrid, and DMA architectures. Moreover, the authors in [21] proposed the location division multiple access (LDMA) where focused beams can be generated to point towards a specific location rather than just a particular direction as the far-field beam steering does, thus increasing the spatial multiplexing gain, and enabling distance level resolution for multi-user communication.
Despite the advantages of near-field transmission, conventional compact antenna arrays usually provide a small near-field region due to the extremely small wavelength at the THz band. Consider a uniform planar array (UPA) with $N$ antennas, the array aperture can be calculated as $S=(\sqrt{N}-1)\sqrt{2}\lambda/2$, where $\lambda$ denotes the wavelength. As the classical boundary between the far-field and the near-field, the Rayleigh distance can then be calculated as $D_{ray}=\frac{2S^{2}}{\lambda}=(\sqrt{N}-1)^{2}\lambda\approx N\lambda$ [22]. Namely, the Rayleigh distance is approximately proportional to the product of the number of antennas and the wavelength. For example, UPA architecture with 1024 antennas at 0.3 THz only provides a near-field region around $D_{ray}\approx 1024\times 0.001=1$m. Therefore, wireless communication usually operates in the far-field under this practical consideration.
Several studies of near-field beamforming assume extremely large number of antenna elements [23, 24, 25] or with uniform linear array (ULA) [26, 27, 28] to increase the near-field region. However, incorporating too many antennas within a limited array space leads to complex hardware configurations, while utilizing ULA may restrict the available spatial dimensions. Instead, a promising beamforming architecture for this problem is to utilize the widely-spaced array (WSA) architecture, which was proposed in our previous work for point-to-point (P2P) communication [14] and also known as sparse array, or line-of-sight (LoS) MIMO [29, 30, 31]. Specifically, the antennas at both the transmitter and the receiver are uniformly divided into $K$ widely-spaced subarrays, i.e., the subarray spacing is larger than half-wavelength. As a result, the effective array aperture can be increased without introducing additional antenna elements, thus enlarging the near-field region while maintaining an acceptable hardware complexity. In multi-user scenarios, however, the available size of user equipment and the number of receiving antennas are often significantly constrained, making it impractical to adopt the WSA. Moreover, a new diagram of cross-field communications needs to be investigated where users residing in both near- and far-field are communicated simultaneously.
To fill this research gap, in this article, we design the hybrid beamforming scheme for multi-user WSA (MU-WSA) where we analyze both the FC and the SC cases. The distinctive contributions of this work are summarized as follows.
  • We analyze the WSA system in multi-user scenarios where the base station is equipped with widely-spaced subarrays such that the near-field region is enlarged without increasing the number of transmit antennas. While at each user, the compact array is adopted to fit the limited size of the user equipment. In this system, a cross-near-and-far-field (CNFF) channel model is established, based on which we analyze the influence of the number of subarrays and the subarray spacing on the sum SE. Specifically, we prove that for a fixed number of subarrays, the channel capacity increases with a larger subarray spacing. Then, based on this result, we propose a search-based algorithm to obtain the optimal design for MU-WSA.
  • We investigate the sub-connected MU-WSA where the analog precoder is shown to have a special block-diagonal form. In this case, we propose a subarray-based alternating optimization (AO) algorithm. Specifically, we decompose the design problem into the analog and digital stages where the block-diagonalization (BD) method is applied in the digital stage to mitigate the interference among users. At the analog stage, we iteratively optimize the precoder matrix of each subarray.
  • We investigate the fully-connected MU-WSA where we propose a low-complexity hybrid beamforming algorithm based on steering vector reconstruction (SVR). Specifically, we apply the far-field steering vector to each subarray in WSA, then reconstruct the vectors based on the CNFF channel model to form a special response vector that is suitable for WSA. This algorithm avoids the utilization of the singular value decomposition (SVD) and can achieve a much lower computational complexity with acceptable performance of the sum SE.
  • We analyze and compare the beam pattern and the sum SE performance of the MU-WSA and the compact array in both 2-D and 3-D cross-field communication systems. We show that the MU-WSA can provide additional distance domain resolution to distinguish the users at the same angle, thus increasing the sum SE by over $60$% at a power of $20$dBm. Furthermore, extensive evaluation has been conducted to numerically investigate the performance of the proposed AO and SVR beamforming algorithms. The results show that the proposed AO algorithm in the SC system can achieve over 80% of the sum (SE) of the FC system while reducing the number of phase shifters by $K^{2}$, thereby lowering power consumption. The SVR algorithm in the FC system can achieve over 95% of the upper bound of SE but takes only 10% of the running time of the singular vector decomposition (SVD)-based algorithms.
The remainder of this paper is organized as follows. In Sec. II, we present the channel model and the system model, based on which we formulate the beamforming design problem. Then, in Sec. III, we analyze the impact of the subarray spacing and the number of subarrays on the spectral efficiency and then design the MU-WSA architecture through a search-based algorithm. Then, in IV, we propose the AO algorithm in the SC system and a low-complexity SVR algorithm in the FC system. Simulation results are shown in Sec. V where we compare the beam pattern of the MU-WSA and compact array architecture in the 2D and 3D scenarios. Then we evaluate the performance of the proposed algorithms. Conclusions are drawn in Sec. VI.
Notations: $\mathbf{A}$ is a matrix, $\mathbf{a}$ is a vector and $a$ is a scalar; $\mathbf{A}(m,n)$ represents the element of $\mathbf{A}$ in row $m$ and column $n$; $\mathbf{I}$ is the identity matrix. $\operatorname{blkdiag}(\mathbf{A}_{1},\cdots,\mathbf{A}_{K})$ is the block diagonal matrix that composed of $\mathbf{A}_{1}$ to $\mathbf{A}_{K}$. $(\cdot)^{T}$ and $(\cdot)^{H}$ represents the transpose and conjugate transpose of a matrix. ||$\cdot$|| ⋅ | and $\|\cdot\|_{F}$ denotes the modulus and the Frobenius norm. $\operatorname{det}(\cdot)$ represents the determinant. $\mathbb{C}^{m\times n}$ is the set of complex-valued matrices of dimension $m\times n$.
Figure 1: Block diagram of the MU-WSA system with hybrid beamforming architecture.
II System and Channel Models
In this section, we introduce the channel model and the system model for the MU-WSA hybrid beamforming architecture.
II-A Cross Near-and-Far-Field (CNFF) Channel Model
As shown in Fig. 1, we consider a downlink MU-WSA system where the BS serves $U$ users simultaneously. At each user, one compact UPA antenna array is equipped due to the limited size of the user equipment. The base station (BS) is equipped with $K$ widely-spaced compact subarrays such that within each subarray, the antennas are spaced with half-wavelength. Because of the small size of the compact antenna array, the Rayleigh distance between the subarray and each user is considered far less than the communication distance $D$, resulting in the far-field propagation scheme. Therefore, according to the planar-wave assumption, the channel matrix between the $u^{\textrm{th}}$ user and the $k^{\textrm{th}}$ subarray at BS can be represented as [7]
$\mathbf{H}^{uk}=\sum_{p=1}^{N_{p}}|\alpha_{p}|e^{-j\frac{2\pi}{\lambda}D_{p}^{uk}}\mathbf{a}_{rp}^{uk}\left(\phi_{rp}^{uk},\theta_{rp}^{uk}\right)\mathbf{a}_{tp}^{uk}\left(\phi_{tp}^{uk},\theta_{tp}^{uk}\right)^{H},$ (1)
where $N_{p}$ is the number of propagation paths for each user, i.e., $N_{p,u}=N_{p},\forall u$. $|\alpha_{p}|$ denotes the modulus of the complex path gain of the $p^{\textrm{th}}$ multipath, and $D_{p}^{uk}$ denotes the transmission distance between the user and the $k^{\textrm{th}}$ subarray. The vectors $\mathbf{a}_{rp}^{uk}\left(\phi_{rp}^{uk},\theta_{rp}^{uk}\right)^{uk}$ and $\mathbf{a}_{tp}^{uk}\left(\phi_{tp}^{uk},\theta_{tp}^{uk}\right)$ are the receive and transmit array response vectors, where $\phi$ and $\theta$ refer to the azimuth and elevation angle, respectively.
Since the widely-spaced subarrays enlarge the effective array aperture of the whole array, the spherical-wave model should be applied among different subarrays. Specifically, by considering different transmission distances and angles of each subarray, the channel matrix of the $u^{\textrm{th}}$ user can be constructed as a combination of the near-field spherical-wave model (SWM) and the far-field PWM, which is written as [32]
$\mathbf{H}_{u}$ $=\begin{bmatrix}\mathbf{H}^{u1}&\cdots&\mathbf{H}^{uk}\end{bmatrix}$ (2)
$=\sum_{p=1}^{N_{p}}|\alpha_{p}|\left[e^{-j\frac{2\pi}{\lambda}D_{p}^{u1}}\mathbf{a}_{rp}^{u1}(\mathbf{a}_{tp}^{u1})^{H},\cdots\right.$
$\quad\quad\quad\quad\quad\quad\left.,\cdots,e^{-j\frac{2\pi}{\lambda}D_{p}^{uK}}\mathbf{a}_{rp}^{uk}(\mathbf{a}_{tp}^{uk})^{H}\right].$
Comparing with conventional PWM where $\operatorname{rank}(\mathbf{H}_{u})=N_{p}$, the upper bound of the rank of $\mathbf{H}_{u}$ increases such that $N_{p}\leq\operatorname{rank}(\mathbf{H}_{u})\leq KN_{p}$, for $K\leq\min\{\frac{N_{t}}{N_{p}},\frac{N_{r}}{N_{p}}\}$.
II-B System Model for MU-WSA
At the receiver side, one compact UPA antenna array with $L_{r}$ RF chains is fully-connected to $N_{r}$ antennas to support the reception of $N_{s}\leq L_{r}$ data streams. While the BS is equipped with $L_{t}$ RF chains that can connect to the $N_{t}$ antennas either through FC or SC. To fully utilize the multiplexing gain of the MIMO system, we assume $UN_{s}=L_{t}=UL_{r}$ in this work.
At the BS, the transmitted symbols $\mathbf{s}=\left[\mathbf{s}_{1}^{T},...,\mathbf{s}_{U}^{T}\right]^{T}$ first go through a diagonal power allocation matrix $\mathbf{P}=\operatorname{blkdiag}\left(\mathbf{P}_{1},...,\mathbf{P}_{U}\right)\in\mathbb{C}^{UN_{s}\times UN_{s}}$, where $\mathbf{P}_{u}$ is the $u^{\textrm{th}}$ diagonal power allocation matrix. Let $\|\mathbf{P}\|^{2}_{F}=P_{t}$, where $P_{t}$ is the total power constraint. After power allocation, the symbols are processed by a baseband precoder $\mathbf{F}_{\mathrm{BB}}\in\mathbb{C}^{L_{t}\times UN_{s}}$, followed by an RF precoder $\mathbf{F}_{\mathrm{RF}}=[\mathbf{F}_{1},\cdots,\mathbf{F}_{K}]$ where $\mathbf{F}_{k}\in\mathbb{C}^{\frac{N_{t}}{K}\times L_{t}}$ represents the analog precoder of each subarray. At the $u^{\textrm{th}}$ user, the received signal is processed by the $u^{\textrm{th}}$ RF combiner $\mathbf{W}_{\mathrm{RF},u}\in\mathbb{C}^{N_{r}\times L_{r}}$, followed by the $u^{\textrm{th}}$ baseband combiner $\mathbf{W}_{\mathrm{BB},u}\in\mathbb{C}^{L_{r}\times N_{s}}$. Therefore the received signal of the $u^{\textrm{th}}$ user is described as
$\tilde{\mathbf{y}}_{u}$ $=\underbrace{\mathbf{W}_{\mathrm{BB},u}^{H}\mathbf{W}_{\mathrm{RF},u}^{H}\mathbf{H}_{u}\mathbf{F}_{\mathrm{RF}}\mathbf{F}_{\mathrm{BB},u}\mathbf{P}_{u}\mathbf{s}_{u}}_{\text{desired signals }}$ (3)
$+\underbrace{\mathbf{W}_{\mathrm{BB},u}^{H}\mathbf{W}_{\mathrm{RF}_{u}}^{H}\mathbf{H}_{u}\sum_{\ell\neq u}\mathbf{F}_{\mathrm{RF}}\mathbf{F}_{\mathrm{BB},{\ell}}\mathbf{P}_{\ell}\mathbf{s}_{\ell}}_{\text{effective interference}}$
$+\underbrace{\mathbf{W}_{\mathrm{BB},u}^{H}\mathbf{W}_{\mathrm{RF},u}^{H}\mathbf{n}_{u}}_{\text{effective noise }},$
where $\mathbf{H}_{u}\in\mathbb{C}^{N_{r}\times N_{t}}$ denotes the channel matrix of the $u^{\textrm{th}}$ user, and $\mathbf{n}_{u}\in\mathbb{C}^{N_{r}\times 1}$ represents the $u^{\textrm{th}}$ noise vector.
Therefore, the achievable spectral efficiency of the $u^{\textrm{th}}$ user is represented as [8]
$SE_{u}=\log_{2}\operatorname{det}\left(\mathbf{I}\right.$ $+\left.\mathbf{R}_{u}^{-1}\mathbf{W}_{\mathrm{BB},{u}}^{H}\mathbf{W}_{\mathrm{RF},{u}}^{H}\mathbf{H}_{u}\mathbf{F}_{\mathrm{RF}}\mathbf{F}_{\mathrm{BB},{u}}\mathbf{P}_{u}\right.$ (4)
×$\left.\times\mathbf{P}_{u}^{H}\mathbf{F}_{\mathrm{BB},{u}}^{H}\mathbf{F}_{\mathrm{RF}}^{H}\mathbf{H}_{u}^{H}\mathbf{W}_{\mathrm{RF},{u}}\mathbf{W}_{\mathrm{BB},{u}}\right),$
where $\mathbf{R}_{u}$ denotes the covariance matrix of the user interference and the noise, i.e.,
$\mathbf{R}_{u}$ $=\sum_{n\neq u}^{U}\mathbf{W}_{\mathrm{BB},{u}}^{H}\mathbf{W}_{\mathrm{RF},{u}}^{H}\mathbf{H}_{u}\mathbf{F}_{\mathrm{RF}}\mathbf{F}_{\mathrm{BB},{n}}\mathbf{P}_{n}$ (5)
×$\times\mathbf{P}_{n}^{H}\mathbf{F}_{\mathrm{BB},{n}}^{H}\mathbf{F}_{\mathrm{RF}}^{H}\mathbf{H}_{u}^{H}\mathbf{W}_{\mathrm{RF},{u}}\mathbf{W}_{\mathrm{BB},{u}}$
$+\sigma_{\mathrm{n}}^{2}\mathbf{W}_{\mathrm{BB},{u}}^{H}\mathbf{W}_{\mathrm{RF},{u}}^{H}\mathbf{W}_{\mathrm{RF},{u}}\mathbf{W}_{\mathrm{BB},{u}}.$
II-B1 Fully-connected MU-WSA
In a fully-connected WSA, each RF chain connects to all the antennas. Therefore, since the analog beamformers $\mathbf{F}_{\mathrm{RF}}$ and $\mathbf{W}_{\mathrm{RF},u}$ are implemented by phase shifters, they should satisfy the constant-modulus constraint, i.e., $\left|\mathbf{F}_{\mathrm{RF}}(m,n)\right|=\frac{1}{\sqrt{N_{t}}}$, and $\left|\mathbf{W}_{\mathrm{RF},{u}}(m,n)\right|=\frac{1}{\sqrt{N_{r}}}$. Additionally, to satisfy the transmitter’s power constraint, we normalize $\mathbf{F}_{\mathrm{BB}}$ such that $\left\|\mathbf{F}_{\mathrm{RF}}\mathbf{F}_{\mathrm{BB}}\right\|_{F}^{2}=UN_{s}$.
II-B2 Sub-connected MU-WSA
In sub-connected WSA, the $L_{t}$ RF chains are divided uniformly into $K$ groups, each group connected to one subarray. Therefore, the number of phase shifters in the sub-connected system can be reduced by a factor of $K^{2}$ compared to the fully-connected system. As a result, the analog precoder has a block-diagonal format as
$\mathbf{F}_{\mathrm{RF}}=\begin{bmatrix}\mathbf{F}_{1}&0&\ldots&0\\ 0&\mathbf{F}_{2}&\ldots&0\\ \ldots&\ldots&\ldots&\ldots\\ 0&\ldots&0&\mathbf{F}_{K}\end{bmatrix},$ (6)
where $\mathbf{F}_{j}\in\mathbb{C}^{\frac{N_{t}}{k}\times\frac{L_{t}}{k}}$ represents the analog precoder for each subarray, and should satisfy $\left|\mathbf{F}_{j}(m,n)\right|=\frac{1}{\sqrt{N_{t}/K}},\forall j$.
II-C Problem Formulation
Based on the system model of the MU-WSA architecture in both FC and SC systems, the beamformers can be designed to maximize the sum SE. Moreover, the subarray spacing $d_{s}$ and the number of subarrays $K$ need to be investigated since they have a significant impact on the system performance [14]. On one hand, as analyzed in Sec. II-A, the rank of the channel matrix is directly related to $K$, thus affecting the SDoF and the sum SE. On the other hand, the values of $K$ and $d_{s}$ affect the position of each antenna, and the channel matrix will change accordingly, resulting in better or worse sub-channels. In practice, due to the space constraint, the size of the BS array aperture must be smaller than a certain threshold, e.g., between $0.2$m and $1.2$[14, 33]. Therefore, the optimization problem aims to jointly design the $(K,d_{s})$ pair along with the hybrid precoders and combiners to maximize the sum SE. Based on (4) and the constraints, the overall optimization problem for FC can be formulated as
$\underset{(K,d_{s}),\mathbf{F}_{\mathrm{RF}},\mathbf{F}_{\mathrm{BB}},\mathbf{W}_{\mathrm{RF}},\mathbf{W}_{\mathrm{BB}},\mathbf{P}}{\operatorname{maximize}}$ $\ SE=\sum_{u=1}^{U}SE_{u}$ (7a)
subject to $\operatorname{Tr}\left(\mathbf{P}^{H}\mathbf{P}\right)=P_{t},$ (7b)
$\left|\mathbf{F}_{\mathrm{RF}}(m,n)\right|^{2}=\frac{1}{N_{t}},\forall m,n,$ (7c)
$\left|\mathbf{W}_{\mathrm{RF},{u}}(m,n)\right|^{2}=\frac{1}{N_{r}},\forall u,m,n,$ (7d)
$S_{t}^{\text{min}}\leq S_{t}\leq S_{t}^{\text{max}},$ (7e)
where constraint (7e) denotes the array aperture constraint where $S_{t}^{\text{min}}$ refers to the case of a compact array. While for SC, the constraint (7c) is replaced by $|\mathbf{F}_{j}(m,n)|^{2}=\frac{1}{N_{t}/K}$. This is a non-convex problem due to the constant-modulus constraints of the analog beamformers. Moreover, the effect of $(K,d_{s})$ is implicit within the expression of the channel matrix, which further complicates the design problem, making it intractable to solve directly. Therefore, we separate this problem into two sub-problems, for which we first design the MU-WSA architecture by analyzing the impact of the number of subarrays and the subarray spacing. Then, we design the hybrid beamforming algorithm to maximize the sum SE.
III Design of MU-WSA Architecture
In this section, we focus on the design of the architecture and neglect the influence of the beamforming algorithm, in order to analyze the upper bound of the sum SE. That is to say, we assume each user transmits its own data streams without interference, such that the performance of each user can be analyzed independently. Considering equal power allocation, the optimal SE, i.e., capacity, for each user is written as
$C_{u}=SE_{u}^{\text{optimal}}=\sum_{i=1}^{\operatorname{rank}(\mathbf{H}_{u})}\log_{2}\left(1+\frac{\rho}{\sigma_{n}^{2}}r_{i}^{2}(\mathbf{H}_{u})\right),$ (8)
where $\rho=P_{t}/\operatorname{rank(\mathbf{H}_{u})}$ denotes the allocated power for each data stream. Since the THz channel has the line-of-sight (LoS)-dominant feature [34], we further simplify the optimization problem as to maximize the capacity of the LoS channel. Moreover, since the correlation of the array response vectors with different elevation angles and different azimuth angles are approximately zero when the number of antennas approaches infinity [21, 35], the rank of the LoS channel approximately equals $K$. To this end, the design problem for the MU-WSA architecture is formed as
$\underset{(K,d_{s})}{\operatorname{maximize}}\ C_{u}^{\text{LoS}}=\sum_{i=1}^{K}\log_{2}\left(1+\frac{\rho}{K\sigma_{n}^{2}}r_{i}^{2}(\mathbf{H}_{u}^{\text{LoS}})\right)$ (9a)
[…]
A gradient-descent method may be feasible to find an optimal solution for this problem in P2P scenarios [14], however, it is impossible to change the array architecture for each user at the same time in multi-user scenarios. To tackle this problem, we present the following Theorem as
Theorem 1
For a given number of subarrays $K$, the capacity of the LoS channel increases monotonically with the subarray distance $d_{s}$ when the transmission distance satisfies $D^{uk}\geq 2\sqrt{2}(\sqrt{N_{r}}-1)S_{t},\forall u,k$.
Proof: We first apply Jensen’s inequality and obtain
$SE^{\text{LoS}}_{\text{optimal}}\leq K\log_{2}(1+\frac{\rho}{K^{2}\sigma_{n}^{2}}\sum_{i=1}^{K}r_{i}^{2}(\mathbf{H}_{u}^{\text{LoS}})).$ (10)
Since $\sum_{i=1}^{K}r_{i}^{2}(\mathbf{H}_{u}^{\text{LoS}})=||\mathbf{H}_{u}^{\text{LoS}}||_{F}^{2}=N_{t}N_{r}\alpha$ is a constant, the maximization of the upper bound becomes to make the inequality holds, namely, the singular values are equal, i.e., $r_{1}=r_{2}=...=r_{K}$. To calculate the singular values of channel $\mathbf{H}_{u}^{\text{LoS}}$, we define $\mathbf{H}_{u}^{\text{LoS}}=\mathbf{A}_{r}\mathbf{\Lambda}(\mathbf{A}_{t})^{H}$ where
$\mathbf{A}_{r}$ $=\begin{bmatrix}\mathbf{a}_{r}^{u1}&\mathbf{a}_{r}^{u2}&\cdots&\mathbf{a}_{r}^{uK}\end{bmatrix},$ (11a)
$\mathbf{A}_{t}$ $=\operatorname{blkdiag}(\mathbf{a}_{t}^{u1},\mathbf{a}_{t}^{u2},\cdots,\mathbf{a}_{t}^{uK})$ (11b)
$=\begin{bmatrix}\bar{\mathbf{a}}_{t}^{u1}&\bar{\mathbf{a}}_{t}^{u2}&\cdots,\bar{\mathbf{a}}_{t}^{uK}\end{bmatrix},$ (11c)
$\mathbf{\Lambda}$ $=\operatorname{blkdiag}(\alpha_{1},\alpha_{2},\cdots,\alpha_{K}).$ (11d)
Therefore, we have
$\mathbf{H}_{u}^{\text{LoS}}(\mathbf{H}_{u}^{\text{LoS}})^{H}$ $=\mathbf{A}_{r}\mathbf{\Lambda}\mathbf{A}_{t}^{H}\mathbf{A}_{t}\mathbf{\Lambda}^{H}\mathbf{A}_{r}^{H}$ (12a)
$=\frac{N_{t}}{K}\mathbf{A}_{r}\mathbf{\Lambda}\mathbf{\Lambda}^{H}\mathbf{A}_{r}^{H}$ (12b)
$=|\alpha|^{2}\frac{N_{t}}{K}\mathbf{A}_{r}\mathbf{A}_{r}^{H}.$ (12c)
Since the elements along the diagonal of $\mathbf{A}_{r}\mathbf{A}_{r}^{H}$ equal $K$, we need to minimize the difference between $\mathbf{A}_{r}\mathbf{A}_{r}^{H}$ and $K\mathbf{I}$. Namely, the optimization problem is equivalent to minimizing the difference between
$f=||\mathbf{A}_{r}\mathbf{A}_{r}^{H}-K\mathbf{I}||_{F}^{2}.$ (13)
The $(i,j)^{th},i\neq j$ element in $\bar{\mathbf{A}}_{r}=\mathbf{A}_{r}\mathbf{A}_{r}^{H}$ can be represented as
$\bar{\mathbf{A}}_{r}(i,j)$ $=\sum_{k=1}^{K}e^{j\frac{2\pi}{\lambda}d_{a}[(j_{1}-i_{1})\sin(\theta_{t}^{uk})\cos(\phi_{t}^{uk})+(j_{2}-i_{2})\sin(\phi_{t}^{uk})]},$ (14a)
where $j_{1},j_{2},i_{1},i_{2}\in[0,\sqrt{Nr}]$ denotes the index on x- and z-axis of the $j^{th}$ and $i^{th}$ antenna. Then, the objective function $f$ can be decomposed as
$f=\sum_{i=1}^{N_{r}}\sum_{j=1}^{N_{r}}|\bar{\mathbf{A}}_{r}(i,j)|^{2}.$ (15)
Since f increases monotonically with $|\bar{\mathbf{A}}_{r}(i,j)|^{2}$, we can further transform the minimization of f as the minimization of function $|\bar{\mathbf{A}}_{r}(i,j)|^{2},\forall i,j$. Define $\bar{\mathbf{A}}_{r}(i,j)=\sum_{k=1}^{K}e^{j\beta_{k,i,j}}$. Since the azimuth angle and the elevation angle can be calculated based on the geometry shown in Fig. 1, i.e.,
$\cos(\phi_{t}^{uk})$ $=\frac{z_{u}-z_{k}}{D^{uk}},$ (16a)
$\sin(\theta_{t}^{uk})$ $=\frac{x_{u}-x_{k}}{D^{uk}\cos{\phi_{t}^{uk}}},$ (16b)
then $\beta_{k,i,j}$ can be transferred into a function of $d_{s}$. Therefore, we have
$g(d_{s})$ $=|\bar{\mathbf{A}}_{r}(i,j)|^{2}$ (17)
$=(\sum_{k=1}^{K}\cos(\beta_{k,i,j}))^{2}+(\sum_{k=1}^{K}\sin(\beta_{k,i,j}))^{2}$
$=K+2\sum_{1\leq k<m\leq K}\cos(\beta_{k,i,j}-\beta_{m,i,j}).$
To prove the monotonicity, we calculate the derivative of $g$ as
$g^{\prime}(d_{s})=2\sum_{k<m}[-\sin(\beta_{k,i,j}-\beta_{m,i,j})\cdot(\frac{\partial(\beta_{k,i,j}-\beta_{m,i,j})}{\partial d_{s}})].$ (18)
Next, we prove that when $d_{s}$ becomes larger, the function $f$ monotonically decreases. This proof is presented in the Appendix. $\blacksquare$
Remark 1: In the proof, we impose a stricter condition than necessary by assuming that each summand $|\mathbf{A}_{r}(i,j)|$ in the function $f$ is monotonically increasing. As a result, the threshold of the transmission distance, i.e., $\tau=2\sqrt{2}(\sqrt{N_{r}}-1)S_{t}^{\text{max}}<D^{uk}$, is in fact a upper bound of the real threshold. That is to say, in practice, the distance can be smaller than $\tau$, while the monotonicity still holds.
Remark 2: Since the threshold $\tau$ is proportional to the square root of the number of the receiving antennas, the theorem is primarily applicable to multi-user scenarios where the user equipment is considered to have much fewer antenna elements than the base station. In contrast, for P2P scenarios where the number of antennas is approximately the same at both ends, the spectral efficiency may exhibit significant fluctuations as $d_{s}$ changes.
Figure 2: Impact of subarray spacing $d_{s}$ and the number of subarrays $k$ on the sum SE.
To directly illustrate the relationship between the channel capacity and the subarray spacing, we present the numerical analysis in Fig. 2. The number of antennas at the BS and each user is set as $N_{t}=1024$ and $N_{r}=64$. Since $K$ denotes the number of subarrays, it must be a factor of the number of antennas. What’s more, for simplicity, we assume $K$ is a square. Therefore, we exhaustively sweep the choices of $K$ in $\mathcal{K}=[1,4,16,64,256]$. For each possible value of $K$, we vary the subarray spacing varies from $1-1000\lambda$. Since the array aperture cannot be infinitely large, we omit the case for larger subarray spacing. In the figure, different colored lines represent different values of $K$, and each line shows the relationship between the capacity and the subarray spacing. At $K=1$, the horizontal line is the baseline representing the compact array case, where the capacity is invariable with the change of the subarray spacing. While for other values of $K$, the channel capacity increases monotonically with the subarray spacing. This is because when $d_{s}$ increases, the difference of the angles $\phi_{p}^{(uk)}$ and $\theta_{p}^{(uk)}$ in (1) becomes larger, so that the difference between the sub-matrices $\mathbf{H}^{(uk)}$ is also larger, leading to a better spatial multiplexing gain. Therefore, based on Theorem 1, and the fact that the choices of $K$ are limited, we propose to first calculate the maximum subarray spacing $d_{s}^{K,\text{max}}$ for each possible value of $K$ under the array aperture threshold, which can be calculated as
$d_{s}^{K,\text{max}}=\frac{\sqrt{2}S_{t}^{\text{max}}+\lambda(\sqrt{N_{t}}-\sqrt{K})}{2(\sqrt{K}-1)}.$ (19)
Then we exhaustively search from all the possible $(K,d_{s})$ pairs to find the solution with the highest channel capacity. This procedure is summarized in Algorithm 1.
Algorithm 1 Search-based design algorithm for subarray sapcing and subarray number in MU-WSA architecture.
0:  Channel matrix $\mathbf{H}$, array aperture threshold $S_{t}^{\text{max}}$.
1:  for $k\in\mathcal{K}$ do
2:     Calculate the maximum available $d_{s}$ through (19) such that the array aperture is no larger than $S_{t}^{\text{max}}$.
3:  end for
4:  Select $(K,d_{s})$ with the highest sum SE.
IV Design of Hybrid Beamforming Algorithm
In this section, we aim to design the hybrid beamformers to maximize the sum SE for all users. To simplify the overall beamforming design problem, we separately design the digital and the analog stages. For the digital beamformers, the BD algorithm is applied to eliminate user interference. For the design of the analog beamformer, two different connection types are considered, i.e., the sub-connected and fully-connected WSA. For the sub-connected case, we propose an alternating-optimization-based algorithm, while for the FC architecture, we propose two low-complexity subarray-based algorithms.
IV-A Digital Precoder and Combiner
For the digital precoder and combiner, we use the baseband block diagonalization method to mitigate the user interference, i.e., the effective interference in (3) should equal to zero. That is, the baseband precoder needs to satisfy $\bar{\mathbf{H}}_{u}\mathbf{F}_{\mathrm{BB},i}=0$ for $i\neq u$, where $\bar{\mathbf{H}}_{u}=\mathbf{W}_{\mathrm{RF},u}^{H}\mathbf{H}_{u}\mathbf{F}_{\mathrm{RF}}$. By defining the interference channel matrix of the $u^{\textrm{th}}$ user as $\hat{\mathbf{H}}_{u}=\begin{bmatrix}\bar{\mathbf{H}}_{1}^{T}&\cdots&\bar{\mathbf{H}}_{u-1}^{T}&\bar{\mathbf{H}}_{u+1}^{T}&\cdots&\bar{\mathbf{H}}_{U}^{T}\end{bmatrix}^{T}$, the baseband precoder should lie in the null space of $\hat{\mathbf{H}}_{u}$. To start with, the singular value decomposition (SVD) of $\hat{\mathbf{H}}_{u}$ is
$\hat{\mathbf{H}}_{u}=\hat{\mathbf{U}}_{u}\hat{\mathbf{\Sigma}}_{u}\left[\hat{\mathbf{V}}_{u}^{(1)},\hat{\mathbf{V}}_{u}^{(0)}\right]^{H},$ (20)
where $\hat{\mathbf{V}}_{u}^{(1)}$ is the first $(U-1)L_{r}$ right singular vectors of $\hat{\mathbf{H}}_{u}$, and $\hat{\mathbf{V}}_{u}^{(0)}$ denotes the orthogonal bases of the null space of $\hat{\mathbf{H}}_{u}$, i.e., $\bar{\mathbf{H}}_{u}\hat{\mathbf{V}}_{i}^{(0)}=0,i\neq u$. As a result, we can eliminate the inter-user interference so that each user can transmit data stream via their own sub-channel $\bar{\mathbf{H}}_{u}\hat{\mathbf{V}}_{u}^{(0)}$. Without the inter-user interference, the sum SE (7a) can then be maximized by achieving the maximum transmission SE of each user. Therefore, we further optimize the precoder and combiner on each sub-channel via SVD, i.e., $\bar{\mathbf{H}}_{u}\hat{\mathbf{V}}_{u}^{(0)}=\mathbf{U}_{u}\mathbf{\Sigma}_{u}\mathbf{V}_{u}^{H}$. We denote $\mathbf{F}_{\mathrm{BB}}=\left[\mathbf{F}_{\mathrm{BB},1},\cdots,\mathbf{F}_{\mathrm{BB},U}\right]$, where $\mathbf{F}_{\mathrm{BB},u}$ is the baseband precoder for the $u^{\textrm{th}}$ data stream. The expressions of the baseband precoder and combiner are given by
$\mathbf{F}_{\mathrm{BB}}$ $=\left[\mathbf{F}_{\mathrm{BB},1},\cdots,\mathbf{F}_{\mathrm{BB},U}\right]$
$=\left[\hat{\mathbf{V}}_{1}^{(0)}\mathbf{V}_{1}^{\left(N_{s}\right)},\cdots,\hat{\mathbf{V}}_{U}^{(0)}\mathbf{V}_{U}^{\left(N_{s}\right)}\right],$ (21a)
$\mathbf{W}_{\mathrm{BB},u}$ $=\mathbf{U}_{u}^{\left(N_{s}\right)},u\in\{1,\ldots,U\}.$ (21b)
where $\mathbf{U}_{u}^{\left(N_{s}\right)}$ and $\mathbf{V}_{u}^{\left(N_{s}\right)}$ are the first $N_{s}$ columns of $\mathbf{U}_{u}$ and $\mathbf{V}_{u}$, respectively.
IV-B Analog Precoder and Combiner
After eliminating the inter-user interference, the sum SE can be written into a more compact form by defining the overall baseband and RF combiner as $\mathbf{W}_{\mathrm{BB}}=\operatorname{blkdiag}\left(\mathbf{W}_{\mathrm{BB},1},\cdots,\mathbf{W}_{\mathrm{BB},U}\right)$ and $\mathbf{W}_{\mathrm{RF}}=\operatorname{blkdiag}\left(\mathbf{W}_{\mathrm{RF},1},\cdots,\mathbf{W}_{\mathrm{RF},U}\right)$. As a result, SE can be expressed as
$SE$ $=\sum_{u=1}^{U}SE_{u}$ (22a)
$\begin{split}&=\log_{2}\operatorname{det}\left(\mathbf{I}+\mathbf{R}^{-1}\mathbf{W}_{\mathrm{BB}}^{H}\mathbf{W}_{\mathrm{RF}}^{H}\mathbf{HF}_{\mathrm{RF}}\mathbf{F}_{\mathrm{BB}}\mathbf{P}\right.\\ &\quad\quad\quad\quad\quad\quad\left.\times\mathbf{P}^{H}\mathbf{F}_{\mathrm{BB}}^{H}\mathbf{F}_{\mathrm{RF}}^{H}\mathbf{H}^{H}\mathbf{W}_{\mathrm{RF}}\mathbf{W}_{\mathrm{BB}}\right)\end{split}$ (22b)
$=\log_{2}\operatorname{det}\left(\mathbf{I}+\frac{\mathbf{W}_{\mathrm{RF}}^{H}\mathbf{HF}_{\mathrm{RF}}\mathbf{F}_{\mathrm{RF}}^{H}\mathbf{H}^{H}\mathbf{W}_{\mathrm{RF}}}{\sigma_{\mathrm{n}}^{2}\mathbf{W}_{\mathrm{RF}}^{H}\mathbf{W}_{\mathrm{RF}}}\right),$ (22c)
As in UM-MIMO systems, the optimal analog combiners typically satisfy $\mathbf{W}_{\mathrm{RF}}^{H}\mathbf{W}_{\mathrm{RF}}\approx\mathbf{I}$ [36], the objective function can be simplified as
$SE\approx\log_{2}\operatorname{det}\left(\mathbf{I}+\sigma_{\mathrm{n}}^{-2}\mathbf{W}_{\mathrm{RF}}^{H}\mathbf{H}\mathbf{F}_{\mathrm{RF}}\mathbf{F}_{\mathrm{RF}}^{H}\mathbf{H}^{H}\mathbf{W}_{\mathrm{RF}}\right).$ (23)
IV-B1 Alternating-Optimization-based Analog Beamforing in Sub-connected MU-WSA
In the sub-connected MU-WSA system where the special block-diagonal structure of the analog precoder should be considered as in (6). To solve this problem, we propose to isolate the contribution of each sub-matrix and optimize one sub-matrix per iteration with other sub-matrices fixed.
We first consider the design of the analog precoder assuming fixed $\mathbf{W}_{\mathrm{RF}}$. Then, the matrix in (23) can be decomposed into four parts, i.e.,
$\log_{2}\operatorname{det}\left(\mathbf{I}+\sigma_{\mathrm{n}}^{-2}\mathbf{F}_{\mathrm{RF}}^{H}\mathbf{A}\mathbf{F}_{\mathrm{RF}}\right)$ (24)
$=$ $\log_{2}\operatorname{det}\left(\begin{array}[]{cc}\mathbf{I}+\frac{\mathbf{F}_{j}^{H}\mathbf{A}_{jj}\mathbf{F}_{j}}{\sigma_{\mathrm{n}}^{2}}&\frac{\mathbf{F}_{j}^{H}\mathbf{A}_{j:}\left(\mathbf{F}_{\mathrm{RF}}\right)_{-j}}{\sigma_{\mathrm{n}}^{2}}\\ \frac{\left(\mathbf{F}_{\mathrm{RF}}\right)_{-j}^{H}\mathbf{A}_{:j}\mathbf{F}_{j}}{\sigma_{\mathrm{n}}^{2}}&\mathbf{I}+\frac{\left(\mathbf{F}_{\mathrm{RF}}\right)_{-j}^{H}(\mathbf{A})_{-jj}\left(\mathbf{F}_{\mathrm{RF}}\right)_{-j}}{\sigma_{\mathrm{n}}^{2}}\end{array}\right)$
=$\overset{(a)}{=}$ $\log_{2}\operatorname{det}\left(\mathbf{B}_{j}\right)+\log_{2}\operatorname{det}\left(\mathbf{I}+\frac{1}{\sigma_{\mathrm{n}}^{2}}\mathbf{F}_{j}^{H}\mathbf{X}_{j}\mathbf{F}_{j}\right),$
where
$\mathbf{A}$ $=\mathbf{H}^{H}\mathbf{W}_{\mathrm{RF}}\mathbf{W}_{\mathrm{RF}}^{H}\mathbf{H}$
$=\begin{bmatrix}\mathbf{A}_{11}&\mathbf{A}_{12}&\ldots&\mathbf{A}_{1k}\\ \mathbf{A}_{21}&\mathbf{A}_{22}&\ldots&\mathbf{A}_{2k}\\ \ldots&\ldots&\ldots&\ldots\\ \mathbf{A}_{k1}&\mathbf{A}_{k2}&\ldots&\mathbf{A}_{kk}\end{bmatrix},\ \mathbf{A}_{ij}\in\mathbb{C}^{\frac{N_{t}}{k}\times\frac{N_{t}}{k}},$ (25a)
$\mathbf{B}_{j}$ $=\mathbf{I}+\frac{\left(\mathbf{F}_{\mathrm{RF}}\right)_{-j}^{H}(\mathbf{A})_{-jj}\left(\mathbf{F}_{\mathrm{RF}}\right)_{-j}}{\sigma_{\mathrm{n}}^{2}},$ (25b)
$\mathbf{X}_{j}$ $=\mathbf{A}_{jj}-\frac{1}{\sigma_{\mathrm{n}}^{2}}\mathbf{A}_{j:}\left(\mathbf{F}_{\mathrm{RF}}\right)_{-j}\mathbf{B}_{j}^{-1}\left(\mathbf{F}_{\mathrm{RF}}\right)_{-j}^{H}\mathbf{A}_{:j},$ (25c)
where $\mathbf{(A)}_{-jj}$ is formed by deleting the rows and columns of $\mathbf{A}$ where the submatrix $\mathbf{A}_{jj}$ is located. $\mathbf{A}_{:j}=\left[\mathbf{A}_{1j}^{T},\cdots,\mathbf{A}_{kj}^{T}\right]^{T}$, and $\mathbf{A}_{j:}=\left[\mathbf{A}_{j1},\cdots,\mathbf{A}_{jk}\right]$. Moreover, $(\mathbf{F}_{\mathrm{RF}})_{-j}$ is the matrix $\mathbf{F}_{\mathrm{RF}}$ except the $j^{\textrm{th}}$ submatrix $\mathbf{F}_{j}$. Step $(a)$ in (24) holds due to the theory that for any block matrix with invertible submatrix $\mathbf{D}$, we have
$\operatorname{det}\left(\begin{bmatrix}\mathbf{A}&\mathbf{B}\\ \mathbf{C}&\mathbf{D}\end{bmatrix}\right)=\operatorname{det}\left(\mathbf{D}\right)\operatorname{det}\left(\mathbf{A}-\mathbf{B}\mathbf{D}^{-1}\mathbf{C}\right).$ (26)
Since $\mathbf{B}_{j}$ in (25b) is not related to $\mathbf{F}_{j}$, we only need to consider the second term in (24). The optimal solution to the unconstrained $\mathbf{F}_{j}$ is then defined by the first $\ell_{t}=L_{t}/k$ columns of the right singular vectors of $\mathbf{X}_{j}=\mathbf{U}_{\mathbf{X}_{j}}\mathbf{\Sigma}_{\mathbf{X}_{j}}\mathbf{V}_{\mathbf{X}_{j}}^{H}$, i.e., $\mathbf{F}_{j}=\mathbf{V}_{\mathbf{X}_{j}}(:,1:\ell_{t})$.
Next, we fix $\mathbf{F}_{\mathrm{RF}}$ and extract the contribution of the analog combiner $\mathbf{W}_{\mathrm{RF}}$ to the sum SE (23). Since the spectral efficiency of each user is only related to its own combiner, we design the analog combiner for each user individually to maximize the overall spectral efficiency, which can be given as
$\log_{2}\operatorname{det}\left(\mathbf{I}+\sigma_{\mathrm{n}}^{-2}\mathbf{W}_{\mathrm{RF}}^{H}\mathbf{D}\mathbf{W}_{\mathrm{RF}}\right)$ (27)
$=$ $\sum_{u=1}^{U}\log_{2}\operatorname{det}\left(\mathbf{I}+\sigma_{\mathrm{n}}^{-2}\mathbf{W}_{\mathrm{RF},u}^{H}\mathbf{D}_{u}\mathbf{W}_{\mathrm{RF},u}\right),$
where $\mathbf{D}=\mathbf{H}^{H}\mathbf{F}_{\mathrm{RF}}\mathbf{F}_{\mathrm{RF}}^{H}\mathbf{H}$ and $\mathbf{D}_{u}=\mathbf{H}_{u}^{H}\mathbf{F}_{\mathrm{RF}}\mathbf{F}_{\mathrm{RF}}^{H}\mathbf{H}_{u}$. Then, the optimal solution of the unconstrained $u^{\textrm{th}}$ analog combiner is given by the first $L_{r}$ columns of the right singular vectors of $\mathbf{D}_{u}=\mathbf{U}_{\mathbf{D}_{u}}\mathbf{\Sigma}_{\mathbf{D}_{u}}\mathbf{V}_{\mathbf{D}_{u}}^{H}$, i.e., $\mathbf{W}_{\mathrm{RF},u}=\mathbf{V}_{\mathbf{D}_{u}}(:,1:L_{r})$.
After finding the unconstrained analog beamformers, we can derive the corresponding solution in the constrained case. According to [10], the elements in the optimal constrained precoder should share the phase of the corresponding element in the unconstrained case. That is, the constrained analog precoder is given by $\mathbf{F}_{j}=\frac{1}{\sqrt{N_{t}}}e^{i\angle{\mathbf{V}_{\mathbf{X}_{j}}(:,1:\ell_{t})}}$. Similarly, the optimal constrained solution for $\mathbf{W}_{\mathrm{RF},u}$ is given by $\mathbf{W}_{\mathrm{RF},u}=\frac{1}{\sqrt{N_{r}}}e^{i\angle{\mathbf{V}_{\mathbf{D}_{u}}(:,1:L_{r})}}$.
A summary of the overall proposed algorithm is presented in Algorithm 2. Specifically, we first optimize the $k$ sub-matrices in $\mathbf{F}_{\mathrm{RF}}$, and then iterate between $\mathbf{F}_{\mathrm{RF}}$ and $\mathbf{W}_{\mathrm{RF}}$ till a stopping criterion triggers. Next, with fixed analog beamforming matrices, we perform a BD method to obtain the baseband precoder and combiner. Finally, we perform the water-filling algorithm for multi-users to obtain the power allocation matrix.
Algorithm 2 Hybrid beamforming with alternating optimization over submatrices in THz multi-user WSA systems
0:  $\mathbf{H}$, $K$
1:  Initialize $\mathbf{W}_{\mathrm{RF}}=\operatorname{blk}\operatorname{dig}\left(\mathbf{W}_{\mathrm{RF},1},\ldots,\mathbf{W}_{\mathrm{RF},U}\right)$ where $\mathbf{W}_{\mathrm{RF},i}\in\mathbb{C}^{N_{r}\times L_{r}}$, and $\mathbf{F}_{\mathrm{RF}}=\left[\mathbf{F}_{1},\ldots,\mathbf{F}_{K}\right]$ where $\mathbf{F}_{\mathrm{i}}=\mathbb{C}^{N_{t}\times\frac{L_{t}}{K}}$.
2:  repeat
3:     Compute $\mathbf{A}=\mathbf{H}^{H}\mathbf{W}_{\mathrm{RF}}\mathbf{W}_{\mathrm{RF}}^{H}\mathbf{H}$
4:     for $j=1$ to $K$ do
5:        Compute $\mathbf{B}_{j}$ through (25b) and $\mathbf{X}_{j}$ through (25c)
6:        $\mathbf{F}_{j}=\mathbf{V}_{\mathbf{X}_{j}}^{(l_{t})}$ is the first $l_{t}$ columns of $\mathbf{V}_{\mathbf{X}_{j}}$
7:     end for
8:     for $u=1$ to $U$ do
9:        $\mathbf{D}_{u}=\mathbf{H}_{u}\mathbf{F}_{\mathrm{RF}}\mathbf{F}_{\mathrm{RF}}^{H}\mathbf{H}_{u}$
10:        $\mathbf{W}_{\mathrm{RF},u}=\mathbf{V}_{\mathbf{D}_{u}}^{(L_{r})}$ where $\mathbf{V}_{\mathbf{D}_{u}}^{(L_{r})}$ is the first $L_{r}$ columns of $\mathbf{V}_{\mathbf{D}_{u}}$
11:     end for
12:  until The stopping criterion trigger is determined by the iteration number N, which will be discussed in the next section.
13:  Compute $\mathbf{F}_{j}=\frac{1}{\sqrt{N_{t}}}e^{i\angle{\mathbf{F}_{j}}}$ and $\mathbf{W}_{\mathrm{RF},u}=\frac{1}{\sqrt{N_{r}}}e^{i\angle{\mathbf{W}_{\mathrm{RF},u}}}$.
14:  Apply the baseband block diagonalization algorithm to calculate $\mathbf{F}_{\mathrm{BB}}$ and $\mathbf{W}_{\mathrm{BB},u}$.
15:  Normalize each column of $\mathbf{F}_{\mathrm{BB}}$ such that $\mathbf{F}_{\mathrm{BB}}(:,i)=\frac{\mathbf{F}_{\mathrm{BB}}(:,i)}{\left\|\mathbf{F}_{\mathrm{RF}}\mathbf{F}_{\mathrm{BB}}(:,i)\right\|_{F}},i\in\left\{1,\ldots,L_{t}\right\}$.
16:  Compute $\mathbf{P}$ based on the water-filling algorithm.
16:  $\mathbf{F}_{\mathrm{RF}},\mathbf{F}_{\mathrm{BB}},(\mathbf{W}_{\mathrm{RF},u},\mathbf{W}_{\mathrm{BB},u})_{u=1:U},\mathbf{P}$
IV-B2 Low-complexity Subarray-based Analog Beamforming in Fully-connected MU-WSA
Now we consider the fully-connected MU-WSA architecture where each subarray is connected to all the RF chains. To reduce the complexity of the analog beamforming, we propose a subarray-based steering vector reconstruction (SVR) scheme, which avoids the utilization of the high-complexity operation of SVD. Since the channel between each subarray and the users can be represented by the far-field PWM, we first simplify the analog beamforming design by designing the analog precoder for each subarray in parallel. Specifically, we aim to maximize the SE for each subarray, i.e.,
$SE_{k}\approx\log_{2}\operatorname{det}\left(\mathbf{I}+\sigma_{\mathrm{n}}^{-2}\mathbf{F}_{\mathrm{RF},k}^{H}\mathbf{H}_{k}^{H}\mathbf{W}_{\mathrm{RF}}\mathbf{W}_{\mathrm{RF}}^{H}\mathbf{H}_{k}\mathbf{F}_{\mathrm{RF},k}\right),$ (28)
where $\mathbf{H}_{k}$ is the PWM between the $k^{\mathrm{th}}$ subarray and all the users, and $\mathbf{F}_{\mathrm{RF},k}$ is the corresponding analog precoding matrix. Assuming a fixed analog combiner, then the analog precoder can be obtained through the beam steering vector. The steering vector between user $u$ and the first subarray can be represented as
$\mathbf{a}_{t}^{u1}=[1,\cdots,e^{j\frac{2\pi}{\lambda}d_{a}((N_{L}-1)\sin\theta_{t}^{u1}\cos(\phi_{t}^{u1})+(N_{W}-1)\sin(\phi_{t}^{u1})}]^{T},$ (29)
where $N_{L}$ and $N_{W}$ are the number of antennas in a subarray at the x-axis and z-axis respectively. Then, the steering vector of the $k^{\mathrm{th}}$ subarray can be constructed by considering the phase difference between different subarrays, i.e.,
$\mathbf{a}_{t}^{uk}$ $=e^{j\frac{2\pi}{\lambda}(D^{uk}-D^{u1})}$ (30)
$\times$ $[1,\cdots,e^{j\frac{2\pi}{\lambda}d_{a}((N_{L}-1)\sin\theta_{t}^{uk}\cos(\phi_{t}^{uk})+(N_{W}-1)\sin(\phi_{t}^{uk})}].$
The distance between the $k^{\mathrm{th}}$ subarray and the user can be represented as
$D^{uk}=\sqrt{(x_{k}-x_{u})^{2}+(y_{k}-y_{u})^{2}+(z_{k}-z_{u})^{2}},$ (31)
where $p_{u}=(x_{u},y_{u},z_{u})$ and $p_{k}=(x_{k},y_{k},z_{k})$ are the position of user $u$ and the $k^{\mathrm{th}}$ subarray, respectively. Denote the distances between different reference antennas on the x- and z-axis as $d_{x}=\frac{(n_{x}-1)}{2}\lambda+d_{s}$ and $d_{z}=\frac{(n_{z}-1)}{2}\lambda+d_{s}$, where $n_{x}=n_{z}=\sqrt{\frac{N_{t}}{K}}$ is the number of antennas in one subarray at the x- and z-axis, and let $p_{1}=(0,0,z_{t})$, then we obtain $p_{k}=(x_{1}+(k_{x}-1)d_{x},0,z_{1}+(k_{z}-1)d_{z})=(x_{1}+\delta d_{x}^{k},0,z_{1}+\delta d_{z}^{k})$, where $k_{x}$ and $k_{z}$ are the x- and z-coordinate of subarray $k$. Therefore, applying the second order Taylor expansion that $\sqrt{1+\psi}\approx 1+\frac{1}{2}\psi-\frac{1}{8}\psi^{2}$, we can derive
$\quad\ =D^{u1}\sqrt{1+\psi}$ (32b)
$\quad\ \approx D^{u1}+\frac{1}{2}\psi^{uk}D^{u1}-\frac{1}{8}(\psi^{uk})^{2}D^{u1},$ (32c)
where
$\psi$ $=\frac{2\delta D_{x}^{k}(x_{1}-x_{u})+2\delta D_{z}^{k}(z_{1}-z_{u})+(\delta D_{x}^{k})^{2}+(\delta D_{z}^{k})^{2}}{(D^{u1})^{2}}$ (33a)
$\begin{split}&=2[k_{z}-1-(k_{x}-1)\sin(\theta_{t}^{u1})]\cos(\phi_{t}^{u1})\frac{d}{D^{u1}}\\ &\quad\quad\quad+[(k_{x}-1)^{2}+(k_{z}-1)^{2}](\frac{d}{D^{u1}})^{2}.\end{split}$ (33b)
(33b) is obtained based on (16), and by considering equal subarray spacing at x- and z-axis, i.e., $d_{x}=d_{z}=d$. Therefore, the distance difference can be calculated as
$\Delta_{D}^{uk}=D^{uk}-D^{u1}=\frac{1}{2}\psi^{uk}D^{u1}-\frac{1}{8}(\psi^{uk})^{2}D^{u1}.$ (34)
As a result, when combining the beam steering vector for each subarray, the overall beam steering vector toward each user is formulated as
$\mathbf{a}_{t}^{(u)}=\begin{bmatrix}\mathbf{a}_{t}^{u1}\\ \vdots\\ \mathbf{a}_{t}^{uk}\end{bmatrix}.$ (35)
Note that when the users are at the far-field region of MU-WSA, such that $D^{uk}\approx D,\forall k$, the reconstructed WSA steering vector is approximately the same as the traditional far-field beam steering vector. Finally, we obtain the resulting analog precoder as
$\mathbf{F}_{\mathrm{RF}}=[\mathbf{a}_{t}^{(1)},\cdots,\mathbf{a}_{t}^{(U)}].$ (36)
The algorithm is summarized in Algorithm 3.
Algorithm 3 Analog beamforming based on steering vector reconstruction (SVR)
0:  Target angle of the first subarray $\{\phi^{u1}\}$, $\{\theta^{u1}\}$, transmission distance $\{D^{u1}\}$
1:  Initialize $\mathbf{F}_{\mathrm{RF}}=[\ ]$.
2:  for $u=1$ to $U$ do
3:     for $k=1$ to $K$ do
4:        Calculate $D^{uk}$ through (32)
5:        Calculate $\theta^{uk}$ and $\phi^{uk}$ through (16).
6:     end for
7:     Construct $\mathbf{a}_{t}^{(u)}$ through (35).
8:     $\mathbf{F}_{\mathrm{RF}}=[\mathbf{F}_{\mathrm{RF}}|\mathbf{a}_{t}^{(u)}]$
9:  end for
9:  $\mathbf{F}_{\mathrm{RF}}$
V Performance Evaluation
In this section, we compare the performance of the MU-WSA and the compact array for both 2-D and 3-D cases. We then evaluate the performance of the proposed algorithm in the MU-WSA architecture and compare it with the state-of-the-art algorithms in the compact architecture under different user scenarios. The simulation setup is summarized in Table I unless otherwise stated.
TABLE I: Simulation Setup.
Parameter Value
Central frequency $f_{c}$ $300$ GHz
Bandwidth B $5$ GHz
Number of multipath components $N_{p}$ 2
Number of transmit antennas $N_{t}$ 1024
Number of receive antennas $N_{r}$ 16
Maximum array aperture $S_{t}^{\text{max}}$ $1$m
2D scenario: BS, UE at height $h_{t}=h_{r}$ $20$m
3D scenario: BS, UE at height $h_{t}$, $h_{r}$ $20$m, $1.5$m
Number of data streams for each user $N_{s}$ 1
Number of users $U$ 20
Number of transmit and receive RF chains $L_{t}$, $L_{r}$ 20, 1
V-A Cross Field Analysis of MU-WSA and Compact Array
In this section, we analyze the beam pattern and the sum spectral efficiency for compact and MU-WSA architecture in the cross near-and-far-field, considering both 2-D and 3-D scenarios.
V-A1 2-D Cross Field Analysis
We first compare the performance of MU-WSA and the compact array in 2-D scenarios where the transmitters and the receivers are at approximately the same height such that only the azimuth angle dominates. In Fig. LABEL:fig:2D-pattern, we analyze the beampattern of MU-WSA and compact array where four users are located at the same azimuth angle, ranging from $1$ to $20$m, represented by the white dot in the figure. Comparing Fig. LABEL:fig_2D:a and LABEL:fig_2D:b, we observe that the far-field analog beam in the compact array fails to distinguish these users since the beamforming gain along the target direction is the same with no distance domain differentiation. While in the MU-WSA, due to the near-field focusing effect, the analog beam pattern depicts the feature of additional distance domain resolution, providing the opportunity to mitigate interference for users with similar angles. Moreover, as the beam is more focused toward the target position rather than a target angle, the power can be more concentrated. To better illustrate this effect, in Fig. LABEL:fig:2D_SE, we show the performance of the sum SE versus the transmission distance. Here we neglect the influence of the beamforming algorithm and consider the upper bound of SE. Fig. LABEL:fig_2D_SE:a shows the sum SE when users are at the same azimuth angle. We observe that with increasing distance between the users and the BS, the MU-WSA architecture first demonstrates a large performance improvement compared to the traditional compact architecture. By contrast, since the compact array fails to cancel the inter-user interference at the same angle, the sum SE drops to approximately zero because of the low SINR. When the communication distance increases to a certain point, the near-field beam-focusing effect vanishes, resulting in a smaller gap between the performance of the MU-WSA and the compact array. On the other hand, for users at different azimuth angles as shown in Fig. LABEL:fig_2D_SE:b, both the MU-WSA and the compact array architecture are able to distinguish the users and the sum SE is almost the same at all distances.
V-A2 3-D Cross Field Analysis
In this section, we consider the 3D scenarios where both the azimuth angle $\phi$ and the elevation angle $\theta$ should be considered. Therefore, even when the users are at the same azimuth angle, the different elevation angles $\phi_{u}$ can support transmission to these users. However, when the distance between users and the BS becomes larger, the elevation angles of different users become approximately the same, i.e., $\phi_{u}\approx\phi,\forall u$. In this case, as shown in Fig. LABEL:fig:3D-pattern, the beamforming gain of the compact and the MU-WSA shows a similar behavior as the 2D case where the MU-WSA provides additional distance domain resolution and depicts the beam focusing effect.
This effect can be further illustrated through the evaluation of the sum SE for users with the same azimuth angles. As shown in Fig. LABEL:fig:3D_SE, we observe that when the transmission distance is small, the performance gap between the MU-WSA and the compact array architecture is relatively small due to the different elevation angles. However, as the communication distance increases, the MU-WSA starts to outperform the compact array. In Fig. LABEL:fig_3D_SE:b where users are all at different azimuth angles, the sum SE is again similar for both architectures.
In Table II, we conclude the MU-WSA and compact array architecture. First, due to the enlarged near-field, MU-WSA has the ability to distinguish users at the same angle while the compact array cannot. Secondly, the MU-WSA architecture takes advantage of the additional spatial degree of freedom such that it is capable of transmitting more data streams than the compact array. However, the MU-WSA will have a relatively large array aperture compared to the compact array and the channel complexity of the cross-field model is higher than the planar-wave model for the compact array architecture.
TABLE II: Comparison of MU-WSA and compact array.
Metric WSA Compact
Channel model CNFF PWM
Array aperture Relatively large Small
Ability to distinguish users at the same angle Yes No
Capable of transmitting number of data streams larger than number of multipath components Yes No
V-B Performance Analysis of Hybrid Beamforming Algorithms
We then analyze the performance of the proposed algorithms in both MU-WSA and the compact array architecture.
In Fig. 11, we first evaluate the sum SE of the proposed algorithms versus the transmit power. The users are distributed randomly within a sector of $120^{\circ}$ angle and a radius ranging from $1$ to $20$m. We compare the performance with the optimal upper bound and an SVD-based benchmark algorithm HySBD-2 proposed in [10]. From the figure, we observe that the MU-WSA architecture always outperforms the compact array regardless of the algorithms used. Specifically, comparing the upper bound of the two architectures, the MU-WSA improves the SE by over $60$% at a power of $20$dBm. In the FC system, the proposed SVR algorithm can achieve over 95% sum SE of the upper bound and the existing SVD-based algorithm, but with much lower computational complexity. Moreover, the proposed AO algorithm in SC achieves over 80% of the sum SE in the FC system with a lower power consumption as the SC system reduces the number of phase shifters by $K^{2}$.
Fig. 12 demonstrates the sum SE versus the number of transmit antennas. We observe that, at both FC and SC systems, when the number of transmit antennas is small, the performance gap between MU-WSA and the compact array is relatively large. As the number of antennas increases, the SE in MU-WSA improves at a more steady rate, while the SE in the compact array increases more rapidly with the growing number of antennas, resulting in a smaller gap of sum SE. This is because when the number of antennas is small, the beamwidth of the compact array is relatively large such that the power cannot be well focused towards a specific direction. When the number of antennas increases, a narrower beam can be generated, resulting in a much higher SE. In contrast, MU-WSA with its fixed array aperture maintains a relatively stable beamwidth and can conduct beam focusing to concentrate the power effectively even with fewer antennas, leading to a more consistent SE improvement. Therefore, the performance of the compact array is more dependent on the number of antennas compared to the MU-WSA.
In Fig. 13, we evaluate the sum SE versus the number of users. We observe that, when the number of users is small, namely, users are sparsely distributed in space, the performance of the WSA and the compact array are similar. However, as the number of users becomes larger to a certain point, the performance of the compact array system experiences a rapid drop due to the similar angles of the users.
In Table III, we analyze the computational complexity and the running time of the proposed algorithms. For SVR which utilizes a predefined steering vector, the precoder at each subarray for each user can be calculated through through simple calculation, resulting in the lowest complexity, i.e., $O(UK)$. While the benchmark algorithm HySBD-2 [10] conducts SVD operation on the channel matrix of all users, resulting in a complexity of $O((UL_{r})^{2}N_{t})$. These two algorithms can be applied to the FC system while the proposed steering vector-based scheme has a much lower complexity. For the more complicated SC beamforming scheme, the proposed AO algorithm is of the highest complexity since it requires the calculation of the inverse of the matrix $\mathbf{B}_{j}\in\mathbb{C}^{\frac{(K-1)L_{t}}{K}\times\frac{(K-1)L_{t}}{K}}$ in (25b), which has a complexity of $O((K-1)^{3}L_{t}^{3}/K^{3})$, and the SVD operation for $\mathbf{X}_{j}\in\mathbb{C}^{\frac{N_{t}}{K}\times\frac{N_{t}}{K}}$, which has a complexity of $O((\frac{N_{t}}{K})^{3})$. Therefore, the overall complexity after $K$ iterations is $O(((K-1)^{3}L_{t}^{3}+N_{t}^{3})/K^{2})$. The running time of the algorithms with MatLab R2022a is also summarized to directly demonstrate the complexity. We observe that the running time corresponds to the complexity analysis, where SVR has the lowest complexity. While AO has the highest complexity but is able to solve the more complex SC problem with acceptable performance loss.
Figure 11: Sum spectral efficiency versus transmit power.
Figure 12: Sum spectral efficiency versus number of transmit antennas.
Figure 13: Sum spectral efficiency versus number of users.
VI Conclusion
TABLE III: Algorithm complexity and running time in MATLAB R2022a.
Algorithm Complexity for analog stage Running time
SVR (FC) $O(UK)$ around 0.02s
HySBD-2 (FC) [10] $O(U^{2}L_{r}^{2}N_{t})$ around 0.18s
AO (SC) $O(((K-1)^{3}L_{t}^{3}+N_{t}^{3})/K^{2})$ around 0.35s
In this paper, we extended the WSA hybrid beamforming architecture into the multi-user scenarios, which can utilize both the angular and distance domain resolutions. Since PWM cannot be applied in the enlarged near-field region of MU-WSA, we investigated the CNFF channel model. Then, by considering both SC and FC systems, we formulated a joint optimization problem to maximize the sum SE. We then decomposed the problem into the design of the MU-WSA architecture, i.e., the subarray spacing $d_{s}$ and the number of subarrays $K$, and the hybrid beamformers. Firstly, based on the analysis of the $(K.d_{s})$ pair, we proposed a search-based algorithm to find the optimal solution for the architecture design. Then, for SC systems where the analog precoder has a block-diagonal format, we propose the AO algorithm to iteratively optimize each submatrix in the analog precoder. In the FC system, a low-complexity SVR algorithm is proposed where the analog beamformers are found through the reconstruction of the far-field steering vectors. We evaluate the beampattern of MU-WSA and the compact array in 2D and 3D cases. We show that in 2D cases, the MU-WSA outperforms the traditional compact array when users are located at the same azimuth angle. While in 3D scenarios, the compact array and the MU-WSA exhibit comparable performance at short distances, but as the distance increases, MU-WSA depicts a gradual superiority over the compact array. When users are located at different angles, the two architectures can achieve similar performance. Therefore, to conclude, the MU-WSA architecture is preferred when the users are more likely to locate at similar angles or directions, while both the MU-WSA and the compact array are feasible when users tend to be scattered in the angle domain. Furthermore, we compare the performance of the proposed AO and SVR algorithms. We show that the proposed AO algorithm in the SC system can achieve over 80% of the sum SE in the FC system but with lower hardware complexity. Moreover, the proposed SVR achieves over 95% of the upper bound of SE, but much lower computational complexity.
Appendix
To prove that $g^{\prime}(d_{s})<0$ when $D^{um}\geq 2\sqrt{2}(\sqrt{N_{r}}-1)S_{t}^{\text{max}},\forall u,m$, we first show $|\beta_{k,i,j}-\beta_{m,i,j}|\leq\pi$. Based on equation (32) and (16), we have
$\sin\theta_{k}\cos\phi_{k}$ $=\frac{x_{u}-x_{k}}{D^{uk}}.$ (37a)
Therefore, $\beta_{k,i,j}$ can be represented as
$\beta_{k,i,j}=\pi\left[(j_{1}-i_{1})\frac{x_{u}-x_{K}}{D^{uk}}+(j_{2}-i_{2})\frac{z_{u}-z_{k}}{D^{uk}}\right],$ (38)
where $(i_{1},j_{1})$ and $(i_{2},j_{2})$ are the x- and z-coordinate of antenna $i$ and $j$, respectively. Define $h(d_{s})=\frac{1}{\pi}(\beta_{k,i,j}-\beta_{m,i,j})$, we next prove $h(d_{s})<1$. Specifically, we can derive
$h(d_{s})=(\beta_{k,i,j}-\beta_{m,i,j})/\pi$ (39)
$=(j_{1}-i_{1})[(x_{u}-x_{k})(\frac{1}{D^{uk}}-\frac{1}{D^{um}})+(m_{x}-k_{x})\frac{d_{s}}{D^{um}}]$
$+(j_{2}-i_{2})[(z_{u}-z_{k})(\frac{1}{D^{uk}}-\frac{1}{D^{um}})+(m_{z}-k_{z})\frac{d_{s}}{D^{um}}]$
$\leq(\sqrt{Nr}-1)[\left|x_{u}-x_{k}\right|\cdot\left|\frac{1}{D^{uk}}-\frac{1}{D^{um}}\right|+(\sqrt{K}-1)\frac{d_{s}}{D^{um}}]$
$+(\sqrt{Nr}-1)[\left|z_{u}-z_{k}\right|\cdot\left|\frac{1}{D^{uk}}-\frac{1}{D^{um}}\right|+(\sqrt{K}-1)\frac{d_{s}}{D^{um}}]$
$=(\sqrt{Nr}-1)\left[\left(|x_{u}-x_{k}|+|z_{u}-z_{k}|\right)\cdot\left|\frac{1}{D^{uk}}-\frac{1}{D^{um}}\right|\right.$
$\quad\quad\quad\quad\quad\quad\left.+2(\sqrt{K}-1)\frac{d_{s}}{D^{um}}\right]$
Since $|D^{uk}-D^{um}|\leq S_{t}$, then by letting $D^{um}=D^{uk}+S_{t}$, we have
$\begin{split}h(d_{s})&\leq(\sqrt{Nr}-1)\left[\frac{(\left|x_{u}-x_{k}\right|+\left|z_{u}-z_{k}\right|)}{D^{uk}}\frac{S_{t}}{D^{uk}+S_{t}}\right.\\ &\quad\quad\quad\quad\quad\quad\left.+2(\sqrt{K}-1)\frac{d_{s}}{D^{um}}\right]\end{split}$ (40a)
$\leq(\sqrt{Nr}-1)(\frac{\sqrt{2}S_{t}}{D^{um}}+2(\sqrt{K}-1)\frac{d_{s}}{D^{um}}),$ (40b)
where (40b) is obtained from the Cauchy–Schwarz inequality that
$|a|+|b|\leq\sqrt{2}\sqrt{a^{2}+b^{2}}\leq\sqrt{a^{2}+b^{2}+c^{2}}.$ (41)
Now, to prove $g(d_{s})<1$, we need to equivalently prove
$\sqrt{2}S_{t}+2(\sqrt{K}-1)d_{s}\leq\frac{D^{um}}{\sqrt{Nr}-1}.$ (42)
Since $\sqrt{2}(\sqrt{K}-1)d_{s}<S_{t}$, we have
$h(d_{s})\leq 2\sqrt{2}S_{t}.$ (43)
Therefore, when the distance between the subarray and the user satisfies $D^{um}\geq 2\sqrt{2}(\sqrt{N_{r}}-1)S_{t},\forall u,m$, the result of $|\beta_{k,i,j}-\beta_{m,i,j}|\leq\pi$ is ensured. Therefore, the original problem is transferred to proving that
$\Delta(d_{s})=2\sum_{k<m}[-(\beta_{k,i,j}-\beta_{m,i,j})\cdot(\frac{\partial(\beta_{k,i,j}-\beta_{m,i,j})}{\partial d_{s}})]<0.$ (44)
Since $\beta_{k,i,j}-\beta_{m,i,j}$ and its derive experience sign consistency, $\Delta(d_{s})<0$ holds. $\blacksquare$